OpenStudy (donnie1999):

HELP PLEASE WILL FAN AND MEDAL!!! A special 8-sided die is marked with the numbers 1 to 8. It is rolled 20 times with these outcomes: 3 4 5 2 7 1 3 7 2 6 2 1 7 3 6 1 8 3 5 6 The experimental probability of rolling a prime number is * BLANK * %, which is * BLANK * % more than the theoretical probability.

OpenStudy (donnie1999):

@welshfella

OpenStudy (mathstudent55):

The probability of rolling a number is the number of way that number can be rolled divided by the total number of outcomes.

OpenStudy (mathstudent55):

There are 8 faces numbered 1 through 8. That means there are 8 different possible outcomes.

OpenStudy (mathstudent55):

How many of those outcomes are prime numbers?

OpenStudy (donnie1999):

@mathstudent55 4 of them are prime numbers.

OpenStudy (mathstudent55):

1, 2, 3, 4, 5, 6, 7, 8 2, 3, 5, 7 are prime You are correct.

OpenStudy (mathstudent55):

There are 4 desired outcomes (2, 3, 5, 7) and there are 8 total possible outcomes. What is the theoretical probability of rolling a prime number?

OpenStudy (donnie1999):

@mathstudent55 would it be 0.5%?

OpenStudy (mathstudent55):

You are on the right track, but you need to be careful with numbers. Theoretical probability: p(rolling a prime number) = 4/8 = 0.5 = 50% 0.5% is incorrect. It's either 50% or 0.5 Ok?

OpenStudy (donnie1999):

yah ok thank

OpenStudy (mathstudent55):

Both 50% and 0.5 probability mean half, which makes sense since half of the total number of outcomes is prime numbers, 4 out of 8.

OpenStudy (mathstudent55):

The theoretical probability of rolling a prime number is 0.5 Now you need the experimental probability. You need to find the probability based on the actual experiment. How many prime numbers were rolled out of a total of how many rolls of the die?

OpenStudy (donnie1999):

12 prime numbers were rolled in the experiment.

OpenStudy (donnie1999):

which is = to 20%

OpenStudy (mathstudent55):

Yes, 12 numbers were prime. How many total rolls of the die?

OpenStudy (donnie1999):

20 rolls

OpenStudy (donnie1999):

do divide 12 by 20 and i got 60%

OpenStudy (mathstudent55):

\(3 ~4 ~5 ~2 ~7 ~1 ~3 ~7 ~2 ~6 ~2 ~1 ~7 ~3 ~6 ~1 ~8 ~3 ~5 ~6\) <--- 20 rolls \(\color{red}{3} ~4 ~\color{red}{5} ~\color{red}{2} ~\color{red}{7} ~1 ~\color{red}{3} ~\color{red}{7} ~\color{red}{2} ~6 ~\color{red}{2} ~1 ~\color{red}{7} ~\color{red}{3} ~6 ~1 ~8 ~\color{red}{3} ~\color{red}{5} ~6\) <--- 12 prime numbers

OpenStudy (mathstudent55):

Correct. Experimental probability: p(roll of a prime) = 12/20 = 0.6 = 60%.

OpenStudy (mathstudent55):

Now the last part. 60% is what percent higher than 50%?

OpenStudy (donnie1999):

10%

OpenStudy (donnie1999):

thanks @mathstudent55

OpenStudy (mathstudent55):

You're welcome.